Linear System
Solve each linear system using the method of substitution.
c) 5x-y=4
3x+y=4
Thank you sooo much

Answer by HyperBrain(694)


[1]5x-y=4
[2]3x+y=4

From [1],
5x-y=4
y=5x-4
Substituting to [2],
3x+5x-4=4
8x=8
x=1
y=5(1)-4=5-4=1
Therefore, x=1 and y=1.
Power up,
HyperBrain!

Help us promote our advocacy:
http://uselessplastic.wordpress.com/
http://uselessplastic.multiply.com/